OpenStudy (anonymous):

I'm solving by using substitution. The Two Equations y=−x/2−4 and −2x−y=−5 the -x is over the 2. Please help. Remember this is solving a system of equations by Substituion

OpenStudy (anonymous):

I have no idea where to start.

OpenStudy (kropot72):

Plug the expression given for y in the first equation into the second equation, then solve the resulting equation in x.

OpenStudy (anonymous):

I know I need to plug it in but I don't know what to do from there. I have a Geometry packet and it says the answer is (6, -7) And when I tried I got (2, -5)

OpenStudy (kropot72):

Plugging in gives: \[-2x-(-\frac{x}{2}-4)=-5\] which simplifies to: \[-2x+\frac{x}{2}+4=-5\ ................(1)\] Can you solve equation (1) for x?

OpenStudy (anonymous):

I don't know how to get rid of the -x/2 and what to multiply 2 by so I can make it just a regular -x

OpenStudy (kropot72):

\[-2x+\frac{x}{2}+4=-5\] Putting the terms in x over a common denominator, we get: \[\frac{-4x+x}{2}+4=-5\] Subtracting 4 from both sides gives us: \[\frac{-4x+x}{2}=-9\ ........(2)\] Can you solve equation (2)?

OpenStudy (anonymous):

Yes Thank you.

OpenStudy (kropot72):

You're welcome :)